OpenStudy (anonymous):

x+2y=13 -2x-3y=-18 solve please?

OpenStudy (mathmate):

x+2y=13 ...(1) -2x-3y=-18 ...(2) 2(1)+(2): y=8 Substitute y=8 in (1) x=13-2*8=-3 So x=-3, y=8